# Discriminating Form and Meaning in Multilingual Models with Minimal-Pair ABX Tasks

### v2

@@ 5.3 Discrimination Scores and Cross-Lingual Transfer
We evaluate cross-lingual transfer on POS, NER, and NLI at the final checkpoint. As originally found by Blevins et al. (2022), transfer accuracy varies widely across source–target pairs (see Appendix H for detailed heatmaps). To test whether ABX discrimination explains this variation, we fit linear regression models predicting transfer accuracy from LD and MD scores between language pairs (see Table 1). We find that LD is a significant negative predictor for both POS and NER. Neither LD nor MD is predictive of NLI performance. This supports the view that strong language-specific encoding can hinder generalization across languages (see Appendix I for visualization). We also test whether ABX language discrimination can guide source language selection for transfer. While it does not consistently identify the single best source, it often selects competitive candidates and outperforms random baselines (see Appendix J).
@@ 6 Conclusion and Future Work
This work introduces ABX-style discrimination metrics for testing how multilingual encoder models discriminate language identity (form) and semantic content (meaning). Adapting ABX to text-based multilingual models, we provide a lightweight, interpretable tool for analysing representational structure.
@@
Future work can build on this in several directions. First, discrimination patterns could be related to typological linguistic features, and it is worth investigating how this typological differences can influences form and content discrimination scores, as previous work has found positive transfer when pairing typologically similar languages Wu and Dredze (2020). Second, while we evaluated tasks spanning syntax and semantics (POS, NER, NLI), deeper semantic tasks could better test the role of meaning discrimination. Third, because ABX metrics are architecture-agnostic, they can be applied to decoder-only LLMs, enabling cross-architecture comparisons. Finally, while ABX does not directly measure representational separation, high discrimination may suggest that a language occupies a distinct subspace. This raises a broader question: how much language sensitivity (i.e., the ability to discriminate languages) can a model have without harming cross-lingual transfer, and how can models balance this trade-off between promoting representational sharing and avoiding interference?
@@
Encoder architecture
Our analysis focuses exclusively on encoder-only architectures, specifically XLM-R. This choice is motivated by the fact that encoder models based on masked language modeling provide stable, structured, layer-wise representations, which are well suited to probing and contrastive analysis. While this makes them a natural starting point for validating our ABX discrimination framework, it remains an open question whether similar dynamics hold for decoder-only or encoder–decoder models, which are trained using autoregressive or sequence-level objectives. Extending our framework to such architectures is an important direction for future work.
@@
Appendix A Languages Used in Evaluations
@@
Appendix B Illustration of Language and Meaning ABX Setups
An example is given below:
  • $X$: “The weather is nice today.” (English ($L_{1}$), meaning $M_{1}$)
  • $A$: “I need to buy groceries.” (English ($L_{1}$), meaning $M_{2}$)
Here is an example for the MD task:
  • $X$: “The weather is nice today.” (English ($L_{1}$), meaning $M_{1}$)
  • $A$: “La météo est bonne aujourd’hui” (French ($L_{2}$), meaning $M_{1}$: “The weather is nice today”)
(a) Language Discrimination
(b) Meaning Discrimination
Appendix C Correlation Analysis Between Language and Meaning Discrimination in XLM-R
@@
We observe a strong overall negative correlation between LD and MD across all language pairs and checkpoints (Spearman’s $\rho=-0.74$, $p<0.001$; Pearson’s $r=-0.68$, $p<0.001$), computed at the (language pair × checkpoint) level. This suggests that, throughout training, language pairs that are more separable in form tend to be less effective in preserving semantic structure.
To verify that this effect is not simply an artifact of training progression, we examine the relationship at the final checkpoint (step 150,000) alone. The inverse correlation persists with even greater magnitude (Spearman’s $\rho=-0.83$, $p<0.001$; Pearson’s $r=-0.72$, $p<0.001$), confirming that the tradeoff between language and meaning discrimination remains pronounced even in the fully trained model. Figure 6 visualizes this relationship: the scatterplot reveals a clear monotonic trend, with almost no high–high co-occurrence (i.e., no language pairs simultaneously scoring high on both MD and LD), which supports the interpretation of a representational tradeoff.
@@
We further analyze the dynamics of this relationship across training by computing correlation coefficients at each checkpoint (Figure 7). Spearman’s correlation remains consistently strong and statistically significant across all training stages, suggesting a stable monotonic inverse relationship. Pearson’s correlation, while also consistently negative, varies in magnitude but remains significant as training progresses, indicating that the relationship is not only ordinal but approximately linear in later stages.
Figure 7: Evolution of Spearman (top) and Pearson (bottom) correlation coefficients between language and meaning discrimination scores across training checkpoints. Statistically significant correlations (p < 0.05) are highlighted.
Appendix D Discrimination scores across languages and checkpoints
To examine how discrimination evolves at the individual language level, we present heatmaps of language and meaning discrimination scores across checkpoints (Figure 8). Scores are normalised per language to highlight relative changes over time. For language discrimination (left), we observe a sharp decline during early training steps for most languages, followed by a partial recovery. However, the timing and extent of this rebound varies across languages, suggesting that some retain language-specific features more robustly. In contrast, meaning discrimination (right) increases steadily for all languages, but again at different rates, with certain languages benefiting earlier from semantically structured representations. These differences may reflect both linguistic factors and data resource disparities. Additional views of final-layer behaviour are included in Figure 9.
@@ Appendix D Discrimination scores across languages and checkpoints
Appendix E Checkpoint-wise Probe Accuracy
@@
Appendix F Additional Probing Analyses
@@
Appendix G ABX-Guided Checkpoint Selection
@@
Appendix H Cross-Lingual Transfer Accuracy Matrices
@@
Appendix I Visualization of Language Discrimination Effects on Cross-Lingual Performance
@@
Appendix J ABX-Guided Source Language Selection
Inspired by our earlier use of ABX scores to guide checkpoint selection (Section G), we investigate whether ABX language discrimination can also inform source language selection in cross-lingual transfer. Specifically, for each target language, we test whether the source language with the lowest ABX language discrimination score yields the highest transfer performance.

### v4

@@ 5.3 Discrimination Scores and Cross-Lingual Transfer
We evaluate cross-lingual transfer on POS, NER, and NLI at the final checkpoint. As originally found by Blevins et al. (2022), transfer accuracy varies widely across source–target pairs (see Appendix I for detailed heatmaps). To test whether ABX discrimination explains this variation, we fit linear regression models predicting transfer accuracy from LD and MD scores between language pairs (see Table 1). We find that LD is a significant negative predictor for both POS and NER. Neither LD nor MD is predictive of NLI performance. This supports the view that strong language-specific encoding can hinder generalization across languages (see Appendix J for visualization). We also test whether ABX language discrimination can guide source language selection for transfer. While it does not consistently identify the single best source, it often selects competitive candidates and outperforms random baselines (see Appendix K).
@@ 6 Conclusion and Future Work
This work introduces ABX-style discrimination metrics for testing how multilingual encoder models discriminate language identity (form) and semantic content (meaning). Adapting ABX to text-based multilingual models, we provide a lightweight, interpretable tool for analyzing representational structure.
@@
Future work can build on this in several directions. First, discrimination patterns could be linked to typological linguistic features, and it is worth investigating how these typological differences can influence form and content discrimination scores, as previous work has found positive transfer when pairing typologically similar languages Wu and Dredze (2020). Second, while we evaluated tasks spanning syntax and semantics (POS, NER, NLI), deeper semantic tasks could better test the role of meaning discrimination. Third, because ABX metrics are architecture-agnostic, they can be applied beyond encoder-only transformers, including decoder-only LLMs and more recent multilingual models, enabling cross-architecture comparisons. Finally, although ABX does not directly measure representational separation, high discrimination may suggest that a language occupies a distinct subspace. This raises a broader question: how much language sensitivity (i.e., the ability to discriminate languages) can a model have without harming cross-lingual transfer, and how can models balance this trade-off between promoting representational sharing and avoiding interference?
@@
Encoder architecture and single-model scope
Our analysis focuses exclusively on encoder-only architectures, specifically XLM-R. This choice is motivated by the fact that encoder models based on masked language modeling provide stable, structured, layer-wise representations, which are well suited to probing and contrastive analysis. We deliberately restrict our experiments to a single model in order to validate the ABX framework itself, rather than provide a broad multi-model benchmark. While this makes them a natural starting point for validating our ABX discrimination framework, it remains an open question whether similar dynamics hold for decoder-only or encoder–decoder models, which are trained using autoregressive or sequence-level objectives. Extending our framework to such architectures and to more recent multilingual models is an important direction for future work.
@@
Appendix A Illustration of Language and Meaning ABX Setups
(a) Language Discrimination
(b) Meaning Discrimination
An example is given below:
  • $X$: “The weather is nice today.” (English ($L_{1}$), meaning $M_{1}$)
  • AA: “I need to buy groceries.” (English ($L_{1}$), meaning $M_{2}$)
Here is an example for the MD task:
  • $X$: “The weather is nice today.” (English ($L_{1}$), meaning $M_{1}$)
  • AA: “La météo est bonne aujourd’hui” (French ($L_{2}$), meaning $M_{1}$: “The weather is nice today”)
Appendix B Pseudocode for ABX Evaluation
We provide pseudocode for the ABX evaluation pipeline. Triplet sampling specifies how $(A,B,X)$ items are constructed for the two tasks: Language Discrimination (LD) and Meaning Discrimination (MD). Scoring then determines whether $X$ is closer to $A$ or $B$ in representation space, using cosine distance at a fixed layer $l$ of the model. The final ABX score is obtained by averaging over many sampled triplets.
Triplet sampling.
Algorithm 1 describes how triplets are drawn for Language Discrimination, and Algorithm 2 for Meaning Discrimination. The corresponding setups are illustrated in Figure 5 (Appendix A), which provides examples of how $A$, $B$, and $X$ are selected in practice.
Remark. In practice, ABX scores are estimated from a large random sample of triplets rather than all possible combinations. The number of triplets controls stability: higher sample sizes give lower variance at the cost of computation. In our experiments, we fix the number of triplets per condition ($\approx$ 100k) to ensure reproducibility and comparability across checkpoints. In a robustness check, we recomputed ABX scores over five different random subsamples (final checkpoint). The average standard deviation was below 0.01, confirming that our results are highly stable with respect to triplet sampling.
Algorithm 1 Sampling ABX triplets for Language Discrimination (LD)
1:Parallel corpus with aligned positions $p$ for languages $\ell_{1},\ell_{2}$
2:for each language pair $(\ell_{1},\ell_{2})$ do
3:  Pick AA in $\ell_{1}$ at position $p_{A}$
4:  Pick $B$ in $\ell_{2}$ at the same position $p_{A}$ // translation of AA; same meaning
5:  Pick $X$ in $\ell_{1}$ at some $p_{X}\neq p_{A}$ // same language as AA; different meaning
6:  Emit $(A,B,X)$
7:  Repeat with $(\ell_{1},\ell_{2})$ swapped
8:  Repeat until desired sample size
9:end for
Algorithm 2 Sampling ABX triplets for Meaning Discrimination (MD)
1:Parallel corpus with aligned positions $p$ for languages $\ell_{1},\ell_{2}$
2:for each language pair $(\ell_{1},\ell_{2})$ do
3:  Pick $X$ in $\ell_{1}$ at position $p_{X}$
4:  Pick AA in $\ell_{2}$ at the same position $p_{X}$ // translation of $X$; same meaning
5:  Pick $B$ in $\ell_{2}$ at some $p_{B}\neq p_{X}$ // same language as AA; different meaning
6:  Emit $(A,B,X)$
7:  Repeat with $(\ell_{1},\ell_{2})$ swapped
8:  Repeat until desired sample size
9:end for
ABX discrimination scoring.
Algorithm 3 shows how a score is computed from each triplet. A score of 1 is assigned if $X$ is closer to $A$ than to $B$, 0 if the opposite, and 0.5 in case of a tie. Averaging across all triplets gives the ABX score for the condition.
Algorithm 3 ABX discrimination scoring
1:Triplets $(A,B,X)$, pretrained model $M$ at checkpoint $c$, tokenizer $T$, layer $l$
2:Define $d(u,v)=1-\frac{u\cdot v}{\lVert u\rVert\lVert v\rVert}$ // cosine distance
3:for each triplet $(A,B,X)$ do
4:  Encode $A,B,X$ to embeddings $e_{A},e_{B},e_{X}$ (mean over subword states at layer $l$)
5:  Compute $d_{AX}\leftarrow d(e_{A},e_{X})$,  $d_{BX}\leftarrow d(e_{B},e_{X})$
6:  $s\leftarrow\begin{cases}1&\text{if }d_{AX}<d_{BX}\\ 0&\text{if }d_{AX}>d_{BX}\\ 0.5&\text{otherwise}\end{cases}$
7:end for
8:return mean ABX score across all triplets
Appendix C Languages Used in Evaluations
@@
Appendix D Correlation Analysis Between Language and Meaning Discrimination in XLM-R
@@
We observe a strong overall negative correlation between LD and MD across all language pairs and checkpoints (Spearman’s $\rho=-0.74$, $p<0.001$; Pearson’s r=0.68r=-0.68, $p<0.001$), computed at the (language pair × checkpoint) level. This suggests that, throughout training, language pairs that are more separable in form tend to be less effective in preserving semantic structure.
To verify that this effect is not simply an artifact of training progression, we examine the relationship at the final checkpoint (step 150,000) alone. The inverse correlation persists with even greater magnitude (Spearman’s $\rho=-0.83$, $p<0.001$; Pearson’s r=0.72r=-0.72, $p<0.001$), confirming that the tradeoff between language and meaning discrimination remains pronounced even in the fully trained model. Figure 6 visualizes this relationship: the scatterplot reveals a clear monotonic trend, with almost no high–high co-occurrence (i.e., no language pairs simultaneously scoring high on both MD and LD), which supports the interpretation of a representational tradeoff.
@@
We further analyze the dynamics of this relationship across training by computing correlation coefficients at each checkpoint (Figure 7). Pearson’s correlation remains consistently negative and statistically significant across all training stages, indicating that the relationship between LD and MD is approximately linear throughout training.
Figure 7: Evolution of Pearson correlation coefficients between language and meaning discrimination scores across training checkpoints. Statistically significant correlations ($p<0.05$) are highlighted.
Appendix E Discrimination scores across languages and checkpoints
To examine how discrimination evolves at the individual language level, we present heatmaps of language and meaning discrimination scores across checkpoints (Figure 8). Scores are normalised per language to highlight relative changes over time. For language discrimination (left), we observe a sharp decline during early training steps for most languages, followed by a partial recovery. However, the timing and extent of this rebound varies across languages, suggesting that some retain language-specific features more robustly. In contrast, meaning discrimination (right) increases steadily for all languages, but again at different rates, with certain languages benefiting earlier from semantically structured representations. These differences may reflect both linguistic factors and data resource disparities. Additional views of final-layer behavior are included in Figure 9.
@@ Appendix E Discrimination scores across languages and checkpoints
Appendix F Checkpoint-wise Probe Accuracy
@@
Appendix G Additional Probing Analyses
@@
Appendix H ABX-Guided Checkpoint Selection
@@
Appendix I Cross-Lingual Transfer Accuracy Matrices
@@
Appendix J Visualization of Language Discrimination Effects on Cross-Lingual Performance
@@
Appendix K ABX-Guided Source Language Selection
Inspired by our earlier use of ABX scores to guide checkpoint selection (Section H), we investigate whether ABX language discrimination can also inform source language selection in cross-lingual transfer. Specifically, for each target language, we test whether the source language with the lowest ABX language discrimination score yields the highest transfer performance.